# IFT-UAM/CSIC-24-112

### v1

@@ I Introduction
Measurements of quantum entanglement between the spins of top quark pairs produced at the Large Hadron Collider (LHC) have recently been performed by the ATLAS ATLAS:2023fsd and CMS CMS:2024pts Collaborations, in a kinematical region near the $t\bar{t}$ production threshold. The ATLAS measurement exhibits a quite sizeable discrepancy with respect to next-to-leading order (NLO) predictions of the Standard Model (SM). On the other hand, the CMS collaboration has found that, remarkably, data is very well explained provided the SM NLO prediction is supplemented by the production of toponium, a yet unobserved $t\bar{t}$ bound state of extremely short lifetime. Bound states of $t\bar{t}$ are a prediction of the SM, and are essentially a non-perturbative phenomenon. In hadron collisions, the formation of a $J^{P}=0^{-}$ colour-singlet resonance is predicted Hagiwara:2008df ; Kiyo:2008bv ; Sumino:2010bv , while in $e^{+}e^{-}$ collisions a $J^{P}=1^{-}$ resonance is expected near threshold Strassler:1990nw ; Sumino:1992ai .
@@
The discovery and characterisation of toponium is, admittedly, a formidable task that requires a very accurate theoretical modeling. As a first step, it is useful to investigate the observables that would allow to investigate its properties, thereby allowing to characterise an excess near threshold as ‘toponium’. This is the purpose of this work. Our goal is not to give precise sensitivity estimations: these would not be realistic in the absence of systematic uncertainties, which can only be evaluated by the experiments. Instead, our focus is on the strategy that experiments could follow, including a comparative study of the observables that might be used for that characterisation. The strategy proposed is, in any case, well motivated by the estimations we provide of the statistical sensitivity to observe deviations.
@@ II Event generation, selection and reconstruction
The effective $gg\eta_{t}$ interaction stands for the triangle loop diagram with a top quark. For our study this is a good approximation because being a spin-zero particle, the only feature required from the production is the cross section, which is fitted from calculations of $t\bar{t}$ production near threshold with non-relativistic effects Hagiwara:2008df ; Kiyo:2008bv ; Sumino:2010bv . The spatial size of toponium aso gives rise to differences in top momentum distrubutions, which can be implemented via a Green function reweighting Fuks:2021xje . However, these differences are rather small, c.f. figs. 13, 14 from Ref. Sumino:2010bv , and not relevant for our analysis.
@@ IV Statistical sensitivity to toponium
Table 1: Expected number of toponium and $b\bar{W}W$ events in selected bins of $m_{t\bar{t}}$, for the baseline benchmark.
@@
For the excess events the expected statistical significance in each bin is computed as $S/\sqrt{B}$. The expected statistical significance of the deviations in the multi-dimensional probability density functions of spin and colour variables is calculated from pseudo-experiments. In each pseudo-experiment, two samples are compared: (i) a sample consisting of $S$ signal and $B$ background events, which is taken as pseudo-data; (ii) a sample of $S+B$ background events. Both signal and random events are randomly chosen from large pools of Monte-Carlo generated events. Repeating the pseudo-experiments $10^{4}$ times, we obtain the probability density function for the expected significance, from which we take the mean value as expected significance. Likewise, one can perform pseudo-experiments comparing two independent samples of $S$ signal and $B$ background events, to verify the level of agreement of pseudo-data with toponium production. As an example, we show the results for spin variables in the bin $m_{t\bar{t}}\leq 360$ GeV in fig. 7. The size of the samples corresponds to Run 2$+$3 luminosity, in the baseline benchmark. The red distribution shows the disagreement of pseudo-data with the $b\bar{b}WW$ hypothesis, while the green distribution exhibits the agreement with the $b\bar{b}WW+\eta_{t}$ hypothesis. Of course, the latter depends on the toponium cross section assumed, but this could be fixed from data by the excess events observed.
@@ Acknowledgements
I gratefully acknowledge many useful discussions with M.L. Mangano and help with toponium modelling, and the CERN Theory Department for hospitality during the realisation of this work. I also thank C. Severi for correspondence regarding Ref. Maltoni:2024csn . This work has been supported by the Spanish Research Agency (Agencia Estatal de Investigación) through projects PID2019-110058GB-C21, PID2022-142545NB-C21 and CEX2020-001007-S funded by MCIN/AEI/10.13039/501100011033, and by Fundação para a Ciência e a Tecnologia (FCT, Portugal) through the project CERN/FIS-PAR/0019/2021.
@@ Appendix A Analysis of the semileptonic decay channel
The expected number of events with $m_{t\bar{t}}\leq 360$ GeV for LHC Run 2 is 95830 for $b\bar{b}WW$ and 4150 for toponium, i.e. around 2.5 times larger than in the dilepton decay mode, c.f. table 1. Note that the ration $S/B=0.041$ is quite close to the one in the dilepton channel — leaving aside the fact that other non-$t\bar{t}$ backgrounds are larger in this channel than in the dilepton mode. Still, the sensitivity to toponium using spin and colour observables is much smaller, $0.8\sigma$ and $1.5\sigma$ respectively. This is quite as expected. For spin observables the use of the hadronic polarimeter, with smaller spin analysing power than the charged lepton, leads to smaller differences in the distributions. In addition, near threshold where the jets are not very energetic, it is found that often one of the two untagged jets used to reconstruct the hadronically-decaying $W$ boson actually does not correspond to the $W$ decay — which spoils the rest-frame angular distributions. (A quality cut on the reconstructed hadronic $W$ mass does not improve the sensitivity.) This smaller sensitivity to toponium is also in agreement with the smaller entanglement significance found by the CMS Collaboration in the semileptonic mode near threshold CMS:2024vqh . For colour observables, the smaller sensitivity can be attributed to the additional hadronic activity that washes out the differences due to the colour connection between the top quarks.

### v2

@@ I Introduction
Measurements of quantum entanglement between the spins of top quark pairs produced at the Large Hadron Collider (LHC) have recently been performed by the ATLAS ATLAS:2023fsd and CMS CMS:2024pts Collaborations, in a kinematical region near the $t\bar{t}$ production threshold. The ATLAS measurement exhibits a quite sizeable discrepancy with respect to next-to-leading order (NLO) predictions of the Standard Model (SM). On the other hand, the CMS collaboration has found that, remarkably, data is very well explained provided the SM NLO prediction is supplemented by the production of toponium, a yet unobserved $t\bar{t}$ bound state of extremely short lifetime. Bound states of $t\bar{t}$ are a prediction of the SM, and are essentially a non-perturbative phenomenon. In hadron collisions, the formation of a $J^{P}=0^{-}$ colour-singlet resonance is predicted Fadin:1990wx ; Hagiwara:2008df ; Kiyo:2008bv ; Sumino:2010bv , while in $e^{+}e^{-}$ collisions a $J^{P}=1^{-}$ resonance is expected near threshold Fadin:1987wz ; Strassler:1990nw ; Sumino:1992ai .
@@
The discovery and characterisation of toponium is, admittedly, a formidable task that requires a very accurate theoretical modeling. As a first step, it is useful to investigate the observables that would reveal its properties, thereby allowing to characterise an excess near threshold as ‘toponium’. This is the purpose of this work. Our goal is not to give precise sensitivity estimations: these would not be realistic in the absence of systematic uncertainties, which can only be evaluated by the experiments. Instead, our focus is on the strategy that experiments could follow, including a comparative study of the observables that might be used for that characterisation. The strategy proposed is, in any case, well motivated by the estimations we provide of the statistical sensitivity to observe deviations.
@@ II Event generation, selection and reconstruction
The effective $gg\eta_{t}$ interaction stands for the triangle loop diagram with a top quark. For our study this is a good approximation because being a spin-zero particle, the only feature required from the production is the cross section, which is fitted from calculations of $t\bar{t}$ production near threshold with non-relativistic effects Hagiwara:2008df ; Kiyo:2008bv ; Sumino:2010bv . The spatial size of toponium also gives rise to differences in top momentum distrubutions, which can be implemented via a Green function reweighting Fuks:2021xje . However, these differences are rather small, c.f. figs. 13, 14 from Ref. Sumino:2010bv , and not relevant for our analysis.
@@ IV Statistical sensitivity to toponium
Table 1: Expected number of toponium and $b\bar{b}WW$ events in selected bins of $m_{t\bar{t}}$, for the baseline benchmark.
@@
For the excess events the expected statistical significance in each bin is computed as $S/\sqrt{B}$. The expected statistical significance of the deviations in the multi-dimensional probability density functions of spin and colour variables is calculated from pseudo-experiments. In each pseudo-experiment, two samples are compared: (i) a sample consisting of $S$ signal and $B$ background events, which is taken as pseudo-data; (ii) a sample of $S+B$ background events. Both signal and random events are randomly chosen from large pools of Monte-Carlo generated events. Repeating the pseudo-experiments $10^{4}$ times, we obtain the probability density function (p.d.f.) for the expected significance, from which we take the mean value as expected significance. Likewise, one can perform pseudo-experiments comparing two independent samples of $S$ signal and $B$ background events, to verify the level of agreement of pseudo-data with toponium production. As an example, we show the results for spin variables in the bin $m_{t\bar{t}}\leq 360$ GeV in fig. 7. The size of the samples corresponds to Run 2$+$3 luminosity, in the baseline benchmark. The red distribution shows the disagreement of pseudo-data with the $b\bar{b}WW$ hypothesis, while the green distribution exhibits the agreement with the $b\bar{b}WW+\eta_{t}$ hypothesis. Of course, the latter depends on the toponium cross section assumed, but this could be fixed from data by the excess events observed.
@@ Acknowledgements
I gratefully acknowledge many useful discussions with M.L. Mangano and help with toponium modelling, and the CERN Theory Department for hospitality during the realisation of this work. I also thank C. Severi for correspondence regarding Ref. Maltoni:2024csn . This work has been supported by the Spanish Research Agency (Agencia Estatal de Investigación) through projects PID2022-142545NB-C21 and CEX2020-001007-S funded by MCIN/AEI/10.13039/501100011033, and by Fundação para a Ciência e a Tecnologia (FCT, Portugal) through the project CERN/FIS-PAR/0019/2021.
@@ Appendix A Analysis of the semileptonic decay channel
The expected number of events with $m_{t\bar{t}}\leq 360$ GeV for LHC Run 2 is 95830 for $b\bar{b}WW$ and 4150 for toponium, i.e. around 2.5 times larger than in the dilepton decay mode, c.f. table 1. Note that the ratio $S/B=0.041$ is quite close to the one in the dilepton channel — leaving aside the fact that other non-$t\bar{t}$ backgrounds are larger in this channel than in the dilepton mode. Still, the sensitivity to toponium using spin and colour observables is much smaller, $0.8\sigma$ and $1.5\sigma$ respectively. This is quite as expected. For spin observables the use of the hadronic polarimeter, with smaller spin analysing power than the charged lepton, leads to smaller differences in the distributions. In addition, near threshold where the jets are not very energetic, it is found that often one of the two untagged jets used to reconstruct the hadronically-decaying $W$ boson actually does not correspond to the $W$ decay — which spoils the rest-frame angular distributions. (A quality cut on the reconstructed hadronic $W$ mass does not improve the sensitivity.) This smaller sensitivity to toponium is also in agreement with the smaller entanglement significance found by the CMS Collaboration in the semileptonic mode near threshold CMS:2024vqh . For colour observables, the smaller sensitivity can be attributed to the additional hadronic activity that washes out the differences due to the colour connection between the top quarks.